OpenStudy (anonymous):

elimination method 0.3x-0.2y=4 0.2x+0.5y= -21/23

OpenStudy (anonymous):

Have a go at this yourself, I will start you off... 0.3x-0.2y=4 [1] 0.2x+0.5y= -21/23 [2] Multiply [1] by 2 and [2] by 3 [1] becomes 0.6 -0.4y=8 call this [3] [2] becomes 0.6 1.5y=-63/23 call this [4] Now your part ..... Subtract line [4] from line [3] to find that x disappears. You can then find y by division. Try it and let me know

OpenStudy (anonymous):

Should read [2] becomes 0.6 + 1.5y=-63/23 call this [4]

OpenStudy (anonymous):

When you have found the value of y from this feed it into any of the two original equations to find x

OpenStudy (anonymous):

I'm lost gianfranco

OpenStudy (anonymous):

Where did youstart to get lost?

OpenStudy (anonymous):

[1] becomes 0.6 -0.4y=8 call this [3] [2] becomes 0.6 + 1.5y=-63/23 call this [4] Subtract [4] from [3] gives ... -1.9x = 8 + 63/23 = 170/23 -1.9x = 170/23 divide by -1.9 on both sides x = 1167/347

OpenStudy (anonymous):

Let us try something else 0.3x-0.2y=4 0.2x+0.5y= -21/23 Multiply everything by 10 3x -2y = 40 [1] 2x + 5y =-210/23 [2] Multiply [1] by 2 and [2] by 3 6x -4y = 80 [1] 6x + 15y =-630/23 [2] Now subtract [2] from [1] -19y = 728/23 divide by -19 = 728/23 now y= 728/23 feed this into any of your equations as the value of y and you will calculate x